OpenStudy (anonymous):

Let f(x) = 2x+3 and g(x) = -x^2 + 5 Find (f of g)(x)

sam (.sam.):

fg(x) f(g(x)) f( -x^2 + 5) 2( -x^2 + 5)+3

OpenStudy (accessdenied):

(f of g)(x) => f(g(x)) Insert x=g(x) for f(x) f(g(x)) = 2*g(x) + 3 = 2(-x^2 + 5) + 3 = -2x^2 + 10 + 3 = -2x^2 + 13

OpenStudy (anonymous):

Thank you both! I really appreciate it.